OpenStudy (anonymous):

what is 2^-2 and its reciprochal

OpenStudy (anonymous):

the reciprocal would be -2^2

OpenStudy (anonymous):

well than I guess I say what is 2^-2

OpenStudy (anonymous):

It's going to be 2^2.

OpenStudy (anonymous):

its 2^2

OpenStudy (anonymous):

\[2^{-2}=\frac{ 1 }{ 2^2}=\frac{ 1 }{ 4 } reciprocal=\frac{ 4 }{ 1 }=4\]